OpenStudy (anonymous):

Find the solutions of the given equation in the indicated interval. 20 (cos(x))^2 + cos(x) − 1 = 0, [0, π]

OpenStudy (anonymous):

what is that 20 doing there?

OpenStudy (anonymous):

its in front of cosx^2

OpenStudy (anonymous):

yeah that one

OpenStudy (anonymous):

yeah its 20 in front of cosx^2...

OpenStudy (anonymous):

then i have absolutely no idea you can see the goofy answers here http://www.wolframalpha.com/input/?i=20+%28cos%28x%29%29^2+%2B+cos%28x%29+%E2%88%92+1+%3D+0%2C

OpenStudy (anonymous):

i am going to make a guess that is it is supposed to be a 2

OpenStudy (anonymous):

\[2(\cos(x))^2+\cos(x)-1=0\] that you can solve

OpenStudy (anonymous):

that one factors as \[(2\cos(x)-1)(\cos(x)+1)=0\] set each factor equal to zero, solve for \(x\)